# Anthropocentric bias in language model evaluation

### v2

@@ 3.2 Mechanistic chauvinism
Mechanistic chauvinism is the tendency to assume that even when LLMs match or exceed average human performance, any substantive difference in strategy counts as evidence that the model’s solution lacks generality. In slogan form: all cognitive kinds are human cognitive kinds. [^3] This bias resembles what Buckner (2013) calls “anthropofabulation”—a combination of (i) human-centric definitions of psychological terms with (ii) an exaggeration of human cognitive abilities. Mechanistic chauvinism differs from anthropofabulation in that it focuses on process rather than semantics or performance. Nevertheless, both forms of bias risk setting standards that no non-human system could meet. Whenever an LLM arrives at a solution through a different computational process than humans use, the mechanistic chauvinist concludes that the LLM lacks genuine competence, regardless of how well it performs.
However, this line of thinking risks obscuring the real capabilities of LLMs. Consider a model that learns a general algorithm for addition, rather than simply memorizing a large set of specific addition problems and their solutions. Now suppose that humans use a somewhat different algorithm for addition. The mere fact that the LLM’s approach differs from the human approach does not mean that the LLM lacks real competence at addition. What matters is whether the LLM has learned a robust, generalizable strategy –– not whether this strategy mirrors human cognition. Indeed, we can easily imagine an LLM outperforming humans at addition while using a distinctly unhuman-like computational process.
The key point is that competence, once abstracted from a narrowly human-centric understanding of the term, should be related to the generality and flexibility of the system’s computations rather than superficial resemblance to the human cognitive architecture. An LLM that relies on a giant lookup table of memorized addition problems would not be competent at addition, because it could not flexibly apply its memorized content-specific transitions to novel cases. But a neural network that learns a general addition algorithm could be seen as competent in the domain, even if the solution it has converged upon is distinctly non-human.
@@ 3.3 An objection
We now anticipate an objection to our discussion of mechanistic chauvinism: since the only indisputable examples of language-driven cognition we have are human, and since the capacities of LLMs are acquired through training on human-produced data, isn’t it appropriate to treat human cognition as the only appropriate yardstick for studying LLMs? Our response is that it depends on how the yardstick is used. While we can – and presumably must – begin our investigation of LLM competencies by comparing them to our own, the role of human cognition as a benchmark can be overstated in two ways. First, the human competencies we use as reference points should be regarded as what philosopher Ali Boyle calls “investigative kinds” (Boyle 2024) - they serve as an initial search template, but not as necessary conditions for cognitive status. Second, questions of the form “Do LLMs have cognitive competence $C$?” should be treated as empirical questions. Though this may seem self-evident, it is a principle that is easy to violate, particularly when we assume that facts about implementation are among the distinguishing features between competence $A$ and competence $B$. If facts about the physical implementation of a competence are included in its definition, then it will be impossible for LLMs, which are implemented in silicon computers, to acquire any of the competencies that humans enjoy. Moreover, this impossibility will be logical, rather than empirical. In order to preserve the empirical character of debates about LLM capacities, therefore, we must focus on the algorithmic level of description, where facts about implementation are explicitly set aside.
@@
The investigation of cognitive capacities in LLMs is best viewed as an iterative, cyclic process in which our conception of the relevant competencies and our understanding of the mechanisms that implement them in LLMs mutually inform and revise each other. Solving the challenge of mapping cognitive tasks to capacities often involves consideration of the underlying mechanisms (Francken et al. 2022). But to home in on the mechanisms responsible for a particular competence, we must make principled decisions about the level of abstraction at which to characterize the mechanism and about how to delineate the boundaries of the mechanism itself. These decisions, in turn, depend on how we conceptualize the cognitive phenomenon we are trying to explain. What results is a investigative process in which our characterization of cognitive tasks, our ontology of capacities, and our understanding of mechanisms evolve in tandem as we search for maximally predictive and explanatory mappings between them. In the case of LLMs, this process may lead us quite far from our initial starting point. Though we inevitably begin by searching for human-like competencies in LLMs, the gradual discovery and refinement of LLM-specific mechanisms may ultimately produce a novel ontology of cognitive kinds, one that is optimized for explaining the distinctive strengths and weaknesses of machine intelligence rather than human intelligence. In this way, the iterative nature of the investigative process allows it to drift away from its anthropocentric origins.
@@ 4 Conclusion
Like humans, LLMs can be right for the wrong reasons; we should not take good performance on various benchmarks, particularly those designed without attention to construct validity and potential confounds, at face value. However, they can also be wrong for the wrong reasons. Auxiliary factors can suppress performance in ways that obscure latent competence, so both successes and failures offer only defeasible evidence. While anthropomorphism has received substantial critical attention, anthropocentric bias has gone largely unexamined. We have argued that it poses a serious obstacle to the objective assessment of LLM capacities, and we offered a taxonomy of its forms to help guide future research.
The most direct way to work out whether auxiliary factors actually cause performance degradation on a task is to look inside the model. As such, mechanistic interpretability techniques are essential for counteracting auxiliary oversight. However, the role of mechanistic interpretability in addressing mechanistic chauvinism is more nuanced. It is possible to decode a feature or circuit of interest from a large neural network even if that feature or circuit does not significantly contribute to the network’s functionality (Makelov et al. 2024, Huang et al. 2023). Consequently, there is a risk of anthropomorphic projection in mechanistic interpretability research. As noted above, we should not combat anthropocentrism by resorting to anthropomorphism. Nevertheless, there exist mechanisms that process information in ways that differ from typical human strategies but are nonetheless robust and general. Mechanistic interpretability research can help identify these mechanisms. When they are discovered, we should not dismiss them simply because they deviate from the human case. Doing so would foreclose the possibility of understanding artificial cognition on its own terms.

### v3

@@ 3.2 Mechanistic chauvinism
Mechanistic chauvinism is the tendency to assume that even when LLMs match or exceed average human performance, any substantive difference in strategy counts as evidence that the model’s solution lacks generality. In slogan form: all cognitive kinds are human cognitive kinds. [^6] This bias resembles what bucknerMorganCanonMeet2013 calls “anthropofabulation”—a combination of (i) human-centric definitions of psychological terms with (ii) an exaggeration of human cognitive abilities. Mechanistic chauvinism differs from anthropofabulation in that it focuses on process rather than semantics or performance. Nevertheless, both forms of bias risk setting standards that no non-human system could meet. Whenever an LLM arrives at a solution through a different computational process than humans use, the mechanistic chauvinist concludes that the LLM lacks genuine competence, regardless of how well it performs.
However, this line of thinking risks obscuring the real capabilities of LLMs. Consider a model that learns a general algorithm for addition, rather than simply memorizing a large set of specific addition problems and their solutions. Now suppose that humans use a somewhat different algorithm for addition. The mere fact that the LLM’s approach differs from the human approach does not mean that the LLM lacks real competence at addition. What matters is whether the LLM has learned a robust, generalizable strategy. Indeed, we can easily imagine an LLM outperforming humans at addition while using a distinctly unhuman-like computational process.This concern is arguably reflected in ongoing debates about the linguistic capacities of LLMs. Current models are not only proficient at generating grammatically well-formed sentences, but can also match human performance on judgments of syntactic acceptability, often aligning closely with human intuitions (huLanguageModelsAlign2024a). Yet, many linguists in the generative tradition, would argue that LLMs lack genuine syntactic competence (chomskyAspectsTheorySyntax1965, chomskyNoamChomskyFalse2023, everaertStructuresNotStrings2015). From this perspective, their success stems not from an internal knowledge of hierarchical syntactic rules, but from their nature as massively scaled statistical models trained to predict the next word. This conclusion, however, is challenged by a growing body of mechanistic evidence suggesting that LLMs do learn and represent hierarchical syntax internally, even if the computations that produce these representations differ from those posited in formal linguistic theories (milliereLanguageModelsModelsforthcoming). This debate thus exemplifies mechanistic chauvinism: if competence is defined by the instantiation of a specific computational architecture hypothesized in humans, then LLMs are deemed incompetent by fiat—thereby dismissing both their behavioral success and the alternative, learned mechanisms that support it.
@@ 3.3 An objection
We now anticipate an objection to our discussion of mechanistic chauvinism: since the only indisputable examples of language-driven cognition are human, and since the capacities of LLMs are acquired through training on human-produced data, isn’t human cognition the most appropriate yardstick for studying LLMs? Our response is that it depends on how the yardstick is used. While we can – and presumably must – begin our investigation of LLM competencies by comparing them to our own, the role of human cognition as a benchmark can be overstated in two ways. First, the human competencies we use as reference points should be regarded as what philosopher Ali Boyle calls “investigative kinds” (boyleDisagreementClassificationComparative2024) - they serve as an initial search template, but not as necessary conditions for cognitive status. Second, questions of the form “Do LLMs have cognitive competence CC?” should be treated as empirical questions. Though this may seem self-evident, it is a principle that is easy to violate, particularly when we assume that facts about implementation are among the distinguishing features between competence $A$ and competence $B$. To preserve the empirical character of debates about the capacities of LLMs, we must focus on the appropriate level of analysis. Following Marr’s classic framework, this means focusing on the computational and algorithmic levels, where we ask both what the system computes and which algorithms enable it to perform that computation. Defining competence at the physical or implementational level—for example, by requiring a biological substrate—would make the question of whether an LLM could possess a human-like competence non-empirical by definition. Therefore, our investigation explicitly sets aside facts about physical implementation.
@@
The investigation of cognitive capacities in LLMs is best viewed as an iterative process in which our conception of LLM competencies, and of the mechanisms that implement them, mutually inform and revise each other. Solving the challenge of mapping cognitive tasks to capacities often involves consideration of the underlying mechanisms (franckenCognitiveOntologySearch2022). But to home in on the mechanisms responsible for a particular competence, we must make principled decisions about the level of abstraction at which to characterize the mechanism and about how to delineate the boundaries of the mechanism itself. These decisions, in turn, depend on how we conceptualize the cognitive phenomenon we are trying to explain. What results is an investigative process in which our characterization of cognitive tasks, our ontology of capacities, and our understanding of mechanisms evolve in tandem. In the case of LLMs, this process may lead us quite far from our initial starting point, and may ultimately produce a novel ontology of cognitive kinds, optimized for explaining the distinctive strengths and weaknesses of machine intelligence rather than human intelligence. In this way, the iterative nature of the investigative process allows it to drift away from its anthropocentric origins.
@@ 4 Conclusion
Like humans, LLMs can be right for the wrong reasons. However, they can also be wrong for the wrong reasons. Auxiliary factors can suppress performance in ways that obscure latent competence. While anthropomorphism has received substantial critical attention, anthropocentric bias has gone largely unexamined. We have argued that it poses a serious obstacle to the objective assessment of LLM capacities, and we offered a taxonomy of its forms to help guide future research.
While carefully designed behavioral experiments are the primary means of identifying performance failures, mechanistic interpretability techniques are uniquely suited to adjudicate between competing hypotheses about the causes of these failures. For instance, when behavioral evidence is ambiguous, intepretability methods can, in principle, help distinguish whether a failure is due to a lack of a competent circuit or due to an auxiliary factor (such as mechanistic interference) suppressing an existing competent circuit. As such, supplementing behavioral with mechanistic evidence is a useful way to counteract auxiliary oversight.
The role of mechanistic interpretability in addressing mechanistic chauvinism is more nuanced. It is possible to decode a feature or circuit of interest from a large neural network even if that feature or circuit does not significantly contribute to the network’s functionality (makelovThisSubspaceYou2024, huangRigorouslyAssessingNatural2023). So there is a risk of anthropomorphic projection in mechanistic interpretability research. Nevertheless, there are examples of robust information processing mechanisms that differ from the ones humans typically use. Mechanistic interpretability research can help identify these mechanisms. When they are discovered, we should not dismiss them simply because they deviate from the human case. Doing so would foreclose the possibility of understanding artificial cognition on its own terms.